ABNORMAL REVERSED SPLITTINGS OF TORSIONAL SUBLEVELS INDUCED BY INVERSION MO-TION IN THE S_{1} STATE OF ACETALDEHYDE

\underline{YUNG-CHING CHOU} and I-CHIA CHEN, Department of Chemistry, National Tsing Hua University, Hsinchu, Taiwan 30013, Republic of China; JON T. HOUGEN, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

The G_{6} group-theoretical high-barrier formalism developed previously for internally rotating and inverting CH_{3}NHD is used to interpret the abnormal torsional splittings in the S_{1} state of acetaldehyde for levels 14^{0−}15^{0}, 14^{0−}15^{1} and14^{0−}15^{2}, in which 14^{0−} denotes the asymmetric inversion tunneling component of the aldehyde hydrogen and 15 denotes vibrational mode for the methyl torsional motion. This formalism, derived using an extended permutation-inversion group G_{6}^{m}, treats simultaneously methyl torsion tunneling, aldehyde hydrogen inversion tunneling and overall rotation. Fits to the rotational states of the four pairs of inversion-torsion vibrational levels (14^{0+}15^{0}, 14^{0−}15^{0}), (14^{0+}15^{1}, 14^{0−}15^{1}), (14^{0+}15^{2}, 14^{0−}15^{2}), and (14^{0+}15^{3}, 14^{0−}15^{3}) are performed, giving rms deviations of 0.003, 0.003, 0.004 and 0.004 cm^{−1}, respectively, in which the deviations are comparable with the experimental uncertainty 0.003 cm^{−1}. For torsional levels lying near the summit of the torsional barrier, this theoretical model including high-order terms provides satisfactory fits to the experimental data. The K structures deviated from a pure torsional rotor are fitted using this formalism and are proved to be from coupling to aldehyde hydrogen inversion.